# Alexander Pugachev1       Alena Fenogenova1,2 Vladislav Mikhailov3       Ekaterina Artemova4

### v1

@@ Abstract
Alexander Pugachev1       Alena Fenogenova1,2 Vladislav Mikhailov3       Ekaterina Artemova4 1Higher School of Economics, 2SaluteDevices 3University of Oslo, 4Toloka AI Correspondence: katya-art@toloka.ai
@@ 1 Introduction
Overall 1003 25.19 $\pm$ 30.12 184.66 $\pm$ 145.72
Table 1: REPA dataset statistics and examples.
@@ 2.2 Response Generation
We generate a response to a query with one of six open-source instruction-finetuned LLMs (see Table 3). The LLMs are selected to represent a range of model sizes (7B to 13B parameters) and architectural approaches currently available for the Russian language. This selection includes both general-purpose LLMs (Llama-2-based) and LLMs specifically fine-tuned for Russian (e.g., ruGPT-3.5-13B). We use the default chat templates and inference hyperparameters provided in the standard configurations by HuggingFace Wolf et al. (2020). Finally, we randomly select responses from two different LLMs for each query, allowing us to compare the responses in a pairwise fashion further.
@@ 2.3 Error Annotation
Each response pair is labeled according to ten potential error types and an aggregated overall criterion. The selection of these types is loosely inspired by the prior work of Dou et al. (2022); Hosking et al. (2024); Ye et al. (2024) and reflects common undesirable behaviors in LLM outputs. The error types are designed to assess LMs from multiple angles. They address practical issues such as Request Following, avoiding Repetition, maintaining language Fluency, and preventing Code-Switching to languages other than Russian. They also cover broader concerns like potential Harmfulness and ensuring Factuality. Additionally, they focus on typical AI-generated text issues—such as logical Contradictions, irrelevant information (Relevance), unexpected Refusal to provide an answer, and Sudden Interruptions. Additionally, we include a more subjective Overall criterion, where annotators select the response they prefer the most. This is most similar to the standard coarse pairwise judgment. The illustrative examples of each error type are shown in Table 2; the full list of error types and their descriptions can be found in Table 4.
@@ 2.4 General Statistics
We summarize the REPA’s general statistics by category, query, and response length in Table 1. Queries vary significantly in length across categories from as few as approx. 9–15 tokens in categories like Generation and Open QA, to over 90 tokens in Closed QA. Responses also vary, with average lengths ranging from approx. 145 tokens in Closed QA to over 209 tokens in Brainstorm.
@@ 3 Experimental Evaluation
Error Type Elo Bradley–Terry Glicko2
Request Following Llama2-13b-chat Llama2-7b-chat Llama2-13b-chat
Factuality Llama2-13b-chat Llama2-7b-chat Llama2-13b-chat
Repetition Llama2-7b-chat Llama2-7b-chat Llama2-7b-chat
Code-Switching Mistral-7B ruGPT-3.5-13B Mistral-7B-Saiga
Relevance Mistral-7B Llama2-7b-chat Mistral-7B-Saiga
Harmfulness Mistral-7B-Saiga Mistral-7B Mistral-7B
Fluency Mistral-7B-Saiga ruGPT-3.5-13B ruGPT-3.5-13B
Contradiction Mistral-7B-Saiga Mistral-7B-Saiga Mistral-7B-Saiga
Sudden Interruption Llama2-13b-chat Llama2-7b-chat Llama2-13b-chat
Refusal Mistral-7B-Saiga ruGPT-3.5-13B ruGPT-3.5-13B
Overall Llama2-13b-chat Llama2-7b-chat Llama2-7b-chat
@@ 3.1 LLM Ranking from Human Preferences
Following ChatBotArena Chiang et al. (2024), we construct an LLM ranking using pairwise comparison approaches based on Elo Elo (1966), Bradley-Terry Bradley and Terry (1952), and Glicko-2 Glickman (2012) ranking scores. The initial Elo rating is set to 1000. The Bradley-Terry algorithm is run for 50 iterations. The parameters $\mu$ and $\phi$ for the Glicko-2 algorithm are set to 1500 and 350, respectively. Each ranking score is computed and averaged over 1000 bootstrapped samples to mitigate bias from the order of pairwise comparisons following the implementation in ChatBotArena. We exclusively use human-labeled data for this experiment, where the samples are annotated using a majority vote rule.
@@ Results
Table 5 presents the model with the highest rank for each error type based on three ranking approaches. Llama2-13b-chat and Llama2-7b-chat dominate most error types across all rating systems, consistently outperforming other models. Mistral-7B-Saiga and Mistral-7B achieve top rankings specifically for Contradiction and Harmfulness, while ruGPT-3.5-13B excels in Fluency and Code-switching. Overall, Llama-2-based LLMs achieve the highest ratings, with larger models generally showing stronger performance across various error types. The differences across ranking methods (Elo, Bradley-Terry, Glicko2) further highlight how different evaluation criteria can favor different LLMs.
@@
Recent studies have discovered several sources of bias that hinder LLM judge performance, including position bias and length bias Zheng et al. (2023); Shi et al. (2024).
@@ 3.3 LLM Ranking from LLM Preferences
Based on experimental results (§3.2), we identify GigaChat Max as the best-performing LLM judge. We use it to rank text generation models, following the setup in §3.1. Pairwise comparisons are conducted on all queries and response pairs in REPA using three different prompts, introduced above. The final rankings are derived from Elo, Bradley-Terry, and Glicko-2 scores.
@@ Results
Error Type Elo Bradley–Terry Glicko2
Request Following Llama2-13b-chat Llama2-13b-chat Llama2-13b-chat
Factuality Llama2-13b-chat Llama2-13b-chat Llama2-13b-chat
Repetition Llama2-13b-chat Llama2-7b-chat Llama2-13b-chat
Code-Switching Llama2-13b-chat Llama2-13b-chat Mistral-7B-Saiga
Relevance Mistral-7B-Saiga Mistral-7B-Saiga Mistral-7B-Saiga
Harmfulness Llama2-13b-chat Llama2-13b-chat Llama2-13b-chat
Fluency Llama2-13b-chat Llama2-13b-chat Llama2-13b-chat
Contradiction Llama2-13b-chat Llama2-13b-chat Llama2-13b-chat
Sudden Interruption Llama2-13b-chat Llama2-13b-chat Llama2-13b-chat
Refusal Llama2-13b-chat Llama2-13b-chat Llama2-13b-chat
Overall Llama2-13b-chat Llama2-13b-chat Llama2-13b-chat
@@ 4 Related Work
Fine-grained Evaluation of Machine-generated Texts
Recent work demonstrates a shift towards nuanced methods for more reliable LLM performance assessment, moving beyond aggregate pairwise judgments. Scarecrow Dou et al. (2022) and TGEA Ge et al. (2022) provide error-annotated datasets for the diagnostic evaluation of generated text, covering a diverse range of linguistic and knowledge-based error types. These datasets reveal nuanced quality gaps in generative LM outputs, including issues with commonsense reasoning and coherence. MISMATCH Murugesan et al. (2023) models human judgments based on 13 fine-grained mismatch error types, building on prior approaches to error detection.
@@
LLM-as-a-Judge Evaluation
Using LLMs as judges has emerged as a scalable and cost-effective alternative to human evaluation for assessing AI model outputs Gu et al. (2025). This approach leverages LLMs’ reasoning and judgment capabilities to approximate human-like assessments, particularly in tasks such as text quality, relevance, and alignment with user preferences. LLMs-as-judges can be categorized into several types, including generic LLM judges Mao et al. (2023); Hackl et al. (2023) (e.g., GPT-4), which are versatile but may lack precision for domain-specific tasks; fine-tuned LLM judges Kim et al. (2024b); Lee et al. (2024); Wang et al. (2023); Zhu et al. (2023) (e.g., PandaLM, Prometheus, Judgelm etc.), which are specifically adapted to evaluation tasks or human preference data for improved accuracy on general or specifics tasks. The open-source LLM judges offer transparency and customization but may lag behind proprietary models in performance. Gureja et al. (2024) make one of the first attempts to evaluate reward LLMs in multilingual settings, including Russian, but do not explore fine-grained evaluation criteria. Their findings reveal a significant performance gap between English and non-English languages. Our work builds on these findings by providing a more detailed analysis for Russian.
Evaluating Russian LMs’
Russian LMs have advanced rapidly, with benchmarks developed to assess their performance in general language understanding Shavrina et al. (2020), zero-shot and few-shot classification Taktasheva et al. (2022), and natural language generation Fenogenova et al. (2024). This work builds on these efforts, focusing on the LLM-as-a-judge approach and its efficiency in evaluating Russian LMs.
@@ 5 Conclusion
This work introduces REPA, one of the first non-English benchmarks for evaluating LLM judge performance according to ten diverse fine-grained criteria. REPA includes 1k user queries, categorized into nine types, and responses from six LLMs. Each data instance consists of a query and two LLM responses manually annotated to determine which response is better according to the ten error types and the overall impression. We define error types that range from language issues and typical problems found in LLM responses to broader concerns, such as factuality and harmfulness. We explore the ability of eight LLM judges to perform similar annotations and their potential limitations, such as position and length biases. Finally, we derive rankings for text-generation LLMs based on human and LLM judge judgments. Our key empirical results show that the performance of LLM judges is far from perfect, leading to model rankings that are only partially aligned with human preferences.

### v2

@@ Abstract
Alexander Pugachev1       Alena Fenogenova1,2 Vladislav Mikhailov3       Ekaterina Artemova4 1HSE University, 2SaluteDevices, 3University of Oslo, 4Toloka AI Correspondence: katya-art@toloka.ai
@@ 1 Introduction
Overall Queries 1003 25.19 $\pm$ 30.12 184.66 $\pm$ 145.72
Data Sources ru_instruct_gpt4: 517 (51.5%), Veles–2.5 337 (33.6%), Tagengo: 121 (12.1%), Aya: 24 (2.4%), Chatbot Arena Conversations: 4 (0.4%)
Table 1: REPA dataset statistics and examples. Data sources distribution is provided for the entire dataset.
@@ 2.2 Response Generation
We generate a response to a query with one of six open-source instruction-finetuned LLMs (see Table 3). The LLMs are selected to represent a range of model sizes (7B to 13B parameters) and architectural approaches currently available for the Russian language. This selection includes both general-purpose LLMs (Llama-2-based) and LLMs specifically fine-tuned for Russian (e.g., ruGPT-3.5-13B). We use the default chat templates and inference hyperparameters provided in the standard configurations by HuggingFace Wolf et al. (2020). Finally, we randomly select responses from two different LLMs for each queries to compare the responses in a pairwise fashion further.
@@ 2.3 Error Annotation
Each response pair is labeled according to ten potential error types and an aggregated overall criterion. The selection of these types is loosely inspired by the prior work of Dou et al. (2022); Hosking et al. (2024); Ye et al. (2024) and reflects common undesirable behaviors in LLM outputs. The error types are designed to assess LMs from multiple angles. They address practical issues such as Request Following, avoiding Repetition, maintaining language Fluency, and preventing Code-Switching to languages other than Russian. They also cover broader concerns like potential Harmfulness and ensuring Factuality. Additionally, they focus on typical AI-generated text issues such as logical Contradictions, irrelevant information (Relevance), unexpected Refusal to provide an answer, and Sudden Interruptions. Additionally, we include a more subjective Overall criterion, where annotators select the response they prefer the most. This is most similar to the standard coarse pairwise judgment. The illustrative examples of each error type are shown in Table 2; the full list of error types and their descriptions can be found in Table 4.
@@ 2.4 General Statistics
We summarize the REPA’s general statistics by category, source, query, and response length in Table 1. Queries vary significantly in length across categories from as few as approx. 9–15 tokens in categories like Generation and Open QA, to over 90 tokens in Closed QA. Responses also vary, with average lengths ranging from approx. 145 tokens in Closed QA to over 209 tokens in Brainstorm.
@@ 3 Experimental Evaluation
Error Type Elo Bradley–Terry Glicko2 Borda Rule
Request Following Llama2-13b-chat Llama2-7b-chat Llama2-13b-chat Llama2-13b-chat
Factuality Llama2-13b-chat Llama2-7b-chat Llama2-13b-chat Llama2-13b-chat
Repetition Llama2-7b-chat Llama2-7b-chat Llama2-7b-chat Llama2-7b-chat
Code-Switching Mistral-7B ruGPT-3.5-13B Mistral-7B-Saiga Mistral-7B-Saiga
Relevance Mistral-7B Llama2-7b-chat Mistral-7B-Saiga Mistral-7B-Saiga
Harmfulness Mistral-7B-Saiga Mistral-7B Mistral-7B Mistral-7B-Saiga
Fluency Mistral-7B-Saiga ruGPT-3.5-13B ruGPT-3.5-13B ruGPT-3.5-13B
Contradiction Mistral-7B-Saiga Mistral-7B-Saiga Mistral-7B-Saiga Mistral-7B-Saiga
Sudden Interruption Llama2-13b-chat Llama2-7b-chat Llama2-13b-chat Llama2-13b-chat
Refusal Mistral-7B-Saiga ruGPT-3.5-13B ruGPT-3.5-13B ruGPT-3.5-13B
Overall Llama2-13b-chat Llama2-7b-chat Llama2-7b-chat Llama2-7b-chat
@@ 3.1 LLM Ranking from Human Preferences
Following ChatBotArena Chiang et al. (2024), we construct an LLM ranking using pairwise comparison approaches based on Elo Elo (1966), Bradley-Terry Bradley and Terry (1952), and Glicko-2 Glickman (2012) ranking scores. The initial Elo rating is set to 1000. The Bradley-Terry algorithm is run for 50 iterations. The parameters $\mu$ and $\phi$ for the Glicko-2 algorithm are set to 1500 and 350, respectively. Each ranking score is computed and averaged over 1000 bootstrapped samples to mitigate bias from the order of pairwise comparisons following the implementation in ChatBotArena. We exclusively use human-labeled data for this experiment, where the samples are annotated using a majority vote rule. We use Borda rule Colombo et al. (2022); Rofin et al. (2023) to aggregate the three rankings obtained.
@@ Results
Table 5 presents the model with the highest rank for each error type based on three ranking approaches. Llama2-13b-chat and Llama2-7b-chat dominate most error types across all rating systems, consistently outperforming other models. Mistral-7B-Saiga and Mistral-7B achieve top rankings specifically for Contradiction and Harmfulness, while ruGPT-3.5-13B excels in Fluency and Code-switching. Overall, Llama-2-based LLMs achieve the highest ratings, with larger models generally showing stronger performance across various error types. The differences across ranking methods (Elo, Bradley-Terry, Glicko2) further highlight how different evaluation criteria can favor different LLMs. The aggregated ranking according to the Borda rule is dominated by Mistral-7B-Saiga in four out of ten error categories, while Llama2-7b-chat is selected as the overall best model.
@@
Recent studies have discovered several sources of bias that hinder LLM judge performance, including position bias and length bias Zheng et al. (2023); Shi et al. (2024) as well as self-preference bias Wataoka et al. (2024).
@@
Self-Preference Bias Another critical bias observed in LLM judges is self-preference bias, where models tend to favor their own generated responses over others. The quantitative analysis of Llama2-13b-chat used as a judge demonstrates this phenomenon with particular clarity: the model selected its own responses in 41.6% of pairwise comparisons, exceeding chance-level expectations. Notably, only 23.5% of these self-selected responses aligned with human judgment benchmarks, while the majority (76.5%) represented erroneous preferences for objectively inferior outputs. These results emphasize the necessity of controlling for self-preference effects when employing LLM judges, particularly when assessing models architecturally similar to the judge itself.
@@ 3.3 LLM Ranking from LLM Preferences
Based on experimental results (§3.2), we identify GigaChat Max as the best-performing LLM judge. We use it to rank text generation models, following the setup in §3.1. Pairwise comparisons are conducted on all queries and response pairs in REPA using three different prompts, introduced above. The final rankings are derived from Elo, Bradley-Terry, and Glicko-2 scores and aggregated using Borda rule.
@@ Results
Error Type Elo Bradley–Terry Glicko2 Borda Rule
Request Following Llama2-13b-chat Llama2-13b-chat Llama2-13b-chat Llama2-13b-chat
Factuality Llama2-13b-chat Llama2-13b-chat Llama2-13b-chat Llama2-13b-chat
Repetition Llama2-13b-chat Llama2-7b-chat Llama2-13b-chat Llama2-13b-chat
Code-Switching Llama2-13b-chat Llama2-13b-chat Mistral-7B-Saiga Llama2-13b-chat
Relevance Mistral-7B-Saiga Mistral-7B-Saiga Mistral-7B-Saiga Mistral-7B-Saiga
Harmfulness Llama2-13b-chat Llama2-13b-chat Llama2-13b-chat Llama2-13b-chat
Fluency Llama2-13b-chat Llama2-13b-chat Llama2-13b-chat Llama2-13b-chat
Contradiction Llama2-13b-chat Llama2-13b-chat Llama2-13b-chat Llama2-13b-chat
Sudden Interruption Llama2-13b-chat Llama2-13b-chat Llama2-13b-chat Llama2-13b-chat
Refusal Llama2-13b-chat Llama2-13b-chat Llama2-13b-chat Llama2-13b-chat
Overall Llama2-13b-chat Llama2-13b-chat Llama2-13b-chat Llama2-13b-chat
@@ 4 Related Work
Fine-grained Evaluation of Machine-generated Texts Recent work demonstrates a shift towards nuanced methods for more reliable LLM performance assessment, moving beyond aggregate pairwise judgments. Scarecrow Dou et al. (2022) and TGEA Ge et al. (2022) provide error-annotated datasets for the diagnostic evaluation of generated text, covering a diverse range of linguistic and knowledge-based error types. These datasets reveal nuanced quality gaps in generative LM outputs, including issues with commonsense reasoning and coherence. MISMATCH Murugesan et al. (2023) models human judgments based on 13 fine-grained mismatch error types, building on prior approaches to error detection.
@@
LLM-as-a-Judge Evaluation Using LLMs as judges has emerged as a scalable and cost-effective alternative to human evaluation for assessing AI model outputs Gu et al. (2025). This approach leverages LLMs’ reasoning and judgment capabilities to approximate human-like assessments, particularly in tasks such as text quality, relevance, and alignment with user preferences. LLMs-as-judges can be categorized into several types, including generic LLM judges Mao et al. (2023); Hackl et al. (2023) (e.g., GPT-4), which are versatile but may lack precision for domain-specific tasks; fine-tuned LLM judges Kim et al. (2024b); Lee et al. (2024); Wang et al. (2023); Zhu et al. (2023) (e.g., PandaLM, Prometheus, Judgelm etc.), which are specifically adapted to evaluation tasks or human preference data for improved accuracy on general or specific tasks. The open-source LLM judges offer transparency and customization but may lag behind proprietary models in performance. Gureja et al. (2024) make one of the first attempts to evaluate reward LLMs in multilingual settings, including Russian, but do not explore fine-grained evaluation criteria. Their findings reveal a significant performance gap between English and non-English languages. Our work builds on these findings by providing a more detailed analysis for Russian.
Evaluating Russian LMs’ Russian LMs have advanced rapidly, with benchmarks developed to assess their performance in general language understanding Shavrina et al. (2020), zero-shot and few-shot classification Taktasheva et al. (2022), and natural language generation Fenogenova et al. (2024). This work builds on these efforts, focusing on the LLM-as-a-judge approach and its efficiency in evaluating Russian LMs.
@@ 5 Conclusion
This work introduces REPA, one of the first non-English benchmarks for evaluating LLM judge performance according to ten diverse fine-grained criteria. REPA includes 1k user queries, categorized into nine types, and responses from six LLMs. Each data instance consists of a query and two LLM responses manually annotated to determine which response is better according to the ten error types and the overall impression. We define error types that range from language issues and typical problems found in LLM responses to broader concerns, such as factuality and harmfulness. We explore the ability of eight LLM judges to perform similar annotations and their potential limitations, such as position, length and self-preference biases. Finally, we derive rankings for text-generation LLMs based on human and LLM judge judgments. Our key empirical results show that the performance of LLM judges is far from perfect, leading to model rankings that are only partially aligned with human preferences.
@@
6 Acknowledgments
This research was supported in part through computational resources of HPC facilities at HSE University Kostenetskiy et al. (2021). This research partially done by A.F. is an output of a research project implemented as part of the Basic Research Program at the National Research University Higher School of Economics (HSE University).